OpenStudy (anonymous):

Deductive reasoning can come in handy for solving puzzles. In this puzzle, each letter represents a different digit. What does the O represent? O + O = E O x O = E M x M = OM

OpenStudy (anonymous):

2 + 2 = 4 2 * 2 = 4 0 * 0 = 2(0) I would say O = 2, E = 4 and M = 0

OpenStudy (anonymous):

hey blexting i dont get these kind of problems either (deducive reasoning) ? is this how you would solve them?

OpenStudy (anonymous):

Sorry, I am not sure what your teacher wants. I just thought about it.